# Attention Is All You Need for LLM-based Code Vulnerability Localization

### v1

@@ Abstract
The rapid expansion of software systems and the growing number of reported vulnerabilities have emphasized the importance of accurately identifying vulnerable code segments. Traditional methods for vulnerability localization, such as manual code audits or rule-based tools, are often time-consuming and limited in scope, typically focusing on specific programming languages or types of vulnerabilities. In recent years, the introduction of large language models (LLMs) such as GPT and LLaMA has opened new possibilities for automating vulnerability detection. However, while LLMs show promise in this area, they face challenges, particularly in maintaining accuracy over longer code contexts. This paper introduces LOVA, a novel framework leveraging the self-attention mechanisms inherent in LLMs to enhance vulnerability localization. Our key insight is that self-attention mechanisms assign varying importance to different parts of the input, making it possible to track how much attention the model focuses on specific lines of code. In the context of vulnerability localization, the hypothesis is that vulnerable lines of code will naturally attract higher attention weights because they have a greater influence on the model’s output. By systematically tracking changes in attention weights and focusing on specific lines of code, LOVA improves the precision of identifying vulnerable lines across various programming languages. Through rigorous experimentation and evaluation, we demonstrate that LOVA significantly outperforms existing LLM-based approaches, achieving up to a 5.3$\times$ improvement in F1-scores. LOVA also demonstrated strong scalability, with up to a 14.6$\times$ improvement in smart contract vulnerability localization across languages like C, Python, Java, and Solidity. Its robustness was proven through consistent performance across different LLM architectures.
1 Introduction
@@
Software vulnerabilities are weaknesses or flaws in a software system that can be exploited by attackers to compromise the security, functionality, or data integrity of the system. As of September 2024, over 240,000 vulnerabilities [5] have been publicly reported in the Common Vulnerabilities and Exposures (CVE) database, and these vulnerabilities can have serious consequences for the security and stability of a software system. Accurately localizing vulnerabilities is increasingly vital and has become a key area of focus [39, 40, 41] in recent years. It allows developers to isolate the specific section or module of code where the flaw exists, enabling targeted and efficient remediation. This precision minimizes the risk of unnecessary changes that can introduce new issues, significantly reducing the time and resources required to fix vulnerabilities. Without precise localization, the process of addressing vulnerabilities can become inefficient, leading to potential system-wide disruptions and further security risks.
@@
Large language models (LLMs), such as GPT [16] and LLaMA [18], are sophisticated AI systems designed to process vast amounts of data to understand and generate text or code with human-like accuracy. These models have become prominent due to their advanced capabilities in tasks such as natural language understanding, translation, summarization, and, more recently, in the domain of software development and security. One emerging use case is vulnerability localization, where LLMs help identify and isolate security flaws within large and complex codebases. Traditionally, vulnerability localization has been a time-consuming and labor-intensive process, often requiring manual code audits or the use of rule-based tools [38, 37]. These methods not only demand significant expertise but are also prone to human error, potentially leading to overlooked vulnerabilities.Additionally, rule-based tools are often limited in scope, typically targeting specific programming languages or certain types of vulnerabilities. LLMs, however, offer a new level of automation and intelligence in this area. LLMs can analyze extensive codebases and detect patterns that are commonly associated with vulnerabilities. These models are trained on massive datasets that include programming languages, software repositories, and vulnerability databases. By pasting the code into an LLM and explicitly requesting it to localize the vulnerability, the LLM can efficiently pinpoint the location of the vulnerable code.
@@
While LLM has demonstrated potential for vulnerability localization, there are several notable limitations. First, LLMs struggle to maintain reasoning accuracy over longer code contexts [11, 12]. For instance, in our exploratory experiments, we observed a significant drop in the output accuracy as the lines of code (LoC) increased: when the LoC exceeded 300, the accuracy fell below 5%. Although more advanced reasoning techniques such as Self-Consistency [3]-—which analyzes the code multiple times using different reasoning paths to ensure consistency—, and rStar [45]-—which uses a self-play generation-discrimination process with Monte Carlo Tree Search (MCTS), where two similar models mutually verify reasoning paths—can improve accuracy, these methods still fail to effectively address the issue of long context degradation. Moreover, they rely on stochastic decoding strategies, which introduce randomness into the results, making it difficult to accurately identify the exact vulnerable lines. Even methods like majority voting resulted in only marginal improvements.
@@
Our work aims to enhance the accuracy of vulnerability localization by leveraging self-attention mechanisms (self-attention allows the model to assign varying importance to different input elements, which is the core of LLMs). This is because self-attention is intrinsically linked to the model’s output and is frequently used to interpret its behavior. We hypothesize that if a LLM has the ability to classify vulnerabilities (e.g., types), it should also inherently possess the capability to pinpoint their exact locations. This is based on the premise that code containing vulnerabilities likely exerts the most substantial impact on the model’s attention mechanisms and its subsequent outputs. However, contemporary large models face challenges in accurately localizing vulnerabilities due to their processing of extensive contextual information. This often results in key lines or code segments being abstracted into more generalized, higher-level information, causing a loss of detail and clarity. Therefore, the critical task is to refine the model’s focus to concentrate specifically on the lines of code that are vulnerable, enhancing the precision of vulnerability detection and localization.
@@
Building on this idea and our exploratory experiments (please refer to §3.2 for validation through a series of experiments), we developed LOVA (LO-cating Vulnerabilities via Attention), which operates in three key stages: code line highlighting, attention calculation, and vulnerability localization. In the code line highlighting stage, two versions of the input are generated: a baseline version and a set of highlighted versions. In the attention calculation stage, attention maps are produced for both versions, and the differences between them are analyzed to assess the impact of highlighting specific lines. Finally, in the vulnerability localization stage, the attention patterns are examined to identify potential vulnerabilities within individual lines of code.
@@
The evaluation of LOVA reveals its superiority in localizing vulnerabilities compared to traditional LLM-based approaches across different programming languages and benchmarks. In terms of effectiveness (RQ1), LOVA achieved a significant improvement in F1-scores on both the Big-Vul and CVEFixes-C datasets, outperforming vanilla output and other sophisticated LLM techniques such as CoT, MoA and rStar by up to 5.3$\times$. This high performance stems from the model’s ability to balance precision and recall, minimizing false positives while maintaining high accuracy in identifying vulnerabilities. Particularly in longer contexts, LOVA demonstrates resilience where other techniques falter, maintaining high Top-N accuracy across various LOC ranges. In terms of scalability (RQ2), LOVA showcases its generalization capabilities by performing effectively across multiple languages, including C, Python, Java, and Solidity. It achieves up to a 14.6$\times$ improvement in contract-level vulnerability localization for smart contracts, underscoring its adaptability to diverse coding environments. The robustness of the system (RQ3) is further demonstrated in its cross-compatibility with various LLM architectures, consistently outperforming baseline methods regardless of the underlying model. Finally, the ablation study (RQ4) highlights the importance of each design component, confirming that the integration of contextual information and attention calculation mechanisms is essential to the method’s overall effectiveness in vulnerability localization.
@@
  • We are the first to discover that the self-attention mechanism can be effectively utilized for vulnerability localization. We demonstrate that by tracking changes in attention weights, it is possible to identify specific lines of code that are likely to contain vulnerabilities.
@@ 2.1 LLM Architecture
A large language model (LLM) is an advanced type of artificial intelligence designed to understand and generate human-like text. These models are typically built using deep learning techniques, specifically neural networks with a large number of parameters, trained on vast amounts of text data. LLM is typically based on the transformer architecture [13].
@@
A transformer is a type of deep learning model architecture. The key innovation of the transformer architecture is its use of self-attention mechanisms, which allow the model to weigh the importance of different words in a sentence, regardless of their position (which will be introduced next). The transformer architecture consists of an encoder and a decoder: the encoder processes the input sequence into context-rich representations, which the decoder then uses, along with previously generated tokens, to produce the final output sequence. Since both the transformer and its sub-components (the encoder and decoder) are capable of handling sequence-to-sequence generation tasks, and each component has different functions, the development of current LLMs has branched into three main technical routes: encoder-decoder (such as T5 [17]), encoder-only (such as BERT [15]), and decoder-only (such as GPT [16] and LLaMA [18]). Due to the advantages of higher training efficiency and better inference performance, state-of-the-art models predominantly use the decoder-only architecture. In this paper, we particularly focus on decoder-only LLMs.
2.2 Tokens and Next Token Prediction
@@
Tokens: Tokens serve as the fundamental building blocks that LLMs use to represent and interpret language. Tokenization is a critical step in preparing text for processing by the model. Before handling raw text input, the LLM employs a tokenizer to break the text or code into smaller units known as tokens. These tokens can represent anything from a single character to a whole word, depending on the complexity of the language and the tokenization method. For example, when a LLM processes code, it breaks down the source code into logical tokens, such as keywords, identifiers, operators, and literals. Each component of the code is treated as a token, which the model processes individually. Similarly, during the output process, the LLM generates tokens instead of producing full text all at once. The tokenizer then converts these tokens back into readable text or code. The use of tokenization enables LLMs to handle different languages and context, regardless of their specific syntax, by converting text or code into tokens that can be systematically processed.
@@
Next Token Prediction: Next token prediction is a core task in training and operating LLM. It involves predicting the next token in a sequence of text based on the preceding tokens. In this process, the model takes an input sequence and generates the most likely next token, continuing this until the desired output length is reached. The training objective of next token prediction can be formalized by the following loss function:
@@
2.3 Self-Attention Mechanism
@@
In decoder-only LLMs, to maintain the autoregressive property during the generation process—where each token is generated based solely on the previously generated tokens without access to subsequent tokens—masks are applied during the calculation of attention weights. This ensures that the attention weights for the current token to any subsequent tokens are set to zero. Furthermore, to enable parallel computation of self-attention, the multi-head attention mechanism divides each embedding into multiple parts, each processed by separate attention heads. The results from these heads are then concatenated and linearly transformed to produce the final output. Consequently, this approach allows the model to simultaneously attend to different aspects of the input sequence, thereby enhancing its ability to capture diverse features.
Attention captures the relevance between tokens, serving as an inherent explainability feature of the transformer architecture. There has been significant research on leveraging attention for the interpretability of language models. Previous research has shown that the self-attention mechanism in LLMs can capture both syntactic and semantic information in code. Specifically, this is achieved by analyzing the attention distribution to identify syntactic or semantic relationships between tokens.
3 Motivation, Key Idea, and Threat Model
@@ 3.1 Motivation
Figure 1: Vulnerable code example
Figure 2: The relationship between Lines of Code (LoC) and Average Top-1 Accuracy
@@
Figure 3: Attention maps for four programs, where (a) represents a vulnerability-free program, (c) shows the program after injecting a vulnerability at line 70 of (a). (b) and (d) represent the programs after highlighting line 70 for (a) and (c), respectively. In each figure, each row corresponds to a line number in the code, while the columns indicate the attention from different decoder layers. Darker colors indicate stronger attention, and the black boxes on the line numbers mark the vulnerable lines identified by the LLM.
While it is true that directly querying LLMs to localize code vulnerabilities, or employing more advanced prompt engineering techniques such as Chain of Thought [1, 2] (i.e., we can apply a step-by-step thought process to systematically break down the potential issues in code) can potentially produce the desired outcome, these methods are constrained by several significant limitations.
First, recent studies [11, 12] have demonstrated that the reasoning capabilities of LLMs deteriorate substantially when working with longer contexts. Since vulnerability localization is inherently a reasoning task that often involves analyzing extended code segments, this degradation poses a significant challenge. For example, for the code shown in Figure 1, when we input the complete file containing the vulnerable code into the LLM for vulnerability localization, only two out of ten attempts successfully identified the vulnerability. Furthermore, as shown in Figure 2, as the lines of code (LoC) increase, the output accuracy of the LLM declines significantly. When the LoC exceeds 300, the accuracy drops to below 5%. While existing advanced reasoning methods (e.g., Self-Consistency [3] and rStar [45]) can improve reasoning to some extent, they are not fundamentally designed to address long context degradation. Consequently, there is still a significant decline in accuracy when dealing with long contexts.
Second, the advanced reasoning methods mentioned earlier typically employ stochastic decoding strategies to explore different reasoning paths, thereby enhancing accuracy. However, this approach introduces randomness into the output, leading to variability that can result in unstable localization outcomes. Consequently, it becomes challenging to pinpoint the exact line containing the vulnerability. Even when methods like majority voting or scoring are used to aggregate multiple outputs, the overall improvement in accuracy is often marginal [4, 10].
3.2 Exploratory Experiments and Key Idea

### v2

# If LLMs Would Just Look: Simple Line-by-line Checking Improves Vulnerability Localization

@@ Abstract.
The rapid expansion of software systems and the growing number of reported vulnerabilities have emphasized the importance of accurately identifying vulnerable code segments. Traditional methods for vulnerability localization, such as manual code audits or rule-based tools, are often time-consuming and limited in scope, typically focusing on specific programming languages or types of vulnerabilities. In recent years, the introduction of large language models (LLMs) such as GPT and LLaMA has opened new possibilities for automating vulnerability localization. However, while LLMs show promise in this area, they face challenges, particularly in maintaining accuracy over longer code contexts. This paper introduces LOVA, a novel framework leveraging the self-attention mechanisms inherent in LLMs to enhance vulnerability localization. Our key insight is that self-attention mechanisms assign varying importance to different parts of the input, making it possible to track how much attention the model focuses on specific lines of code. In the context of vulnerability localization, the hypothesis is that vulnerable lines of code will naturally attract higher attention weights because they have a greater influence on the model’s output. By systematically tracking changes in attention weights and focusing on specific lines of code, LOVA improves the precision of identifying vulnerable lines across various programming languages. Through rigorous experimentation and evaluation, we demonstrate that LOVA significantly outperforms existing LLM-based approaches, achieving up to a 5.3$\times$ improvement in F1-scores. LOVA also demonstrated strong scalability, with up to a 14.6$\times$ improvement in smart contract vulnerability localization across languages like C, Python, Java, and Solidity. Its robustness was proven through consistent performance across different LLM architectures.
Large language models, Vulnerability localization
ccs: Security and privacy Software security engineeringccs: Computing methodologies Machine learning
1. Introduction
@@
Software vulnerabilities are weaknesses or flaws that attackers can exploit to compromise a system’s security, functionality, or data integrity. As of September 2024, over 240,000 vulnerabilities (MITRE, 2024) have been publicly reported in the Common Vulnerabilities and Exposures (CVE) database, and these vulnerabilities can have serious consequences for the security and stability of a software system. Accurately localizing vulnerabilities is increasingly vital and has become a key area of focus (Abreu et al., 2007; Moon et al., 2014; Li et al., 2019) in recent years. It allows developers to isolate the specific section or module of code where the flaw exists, enabling targeted and efficient remediation. This precision minimizes the risk of unnecessary changes that can introduce new issues, significantly reducing the time and resources required to fix vulnerabilities. Without precise localization, the process of addressing vulnerabilities can become inefficient, leading to potential system-wide disruptions and further security risks.
@@
Large language models (LLMs), such as GPT (Achiam et al., 2023) and LLaMA (Touvron et al., 2023), are sophisticated AI systems designed to process vast amounts of data to understand and generate text or code with human-like accuracy. These models have become prominent due to their advanced capabilities in tasks such as natural language understanding, translation, summarization, and, more recently, in the domain of software development and security. One emerging use case is vulnerability localization, where LLMs help identify and isolate security flaws within large and complex codebases. Traditionally, vulnerability localization has been a time-consuming and labor-intensive process, often requiring manual code audits or the use of rule-based tools (Wong et al., 2016; Zou et al., 2021). These methods not only demand significant expertise but are also prone to human error, potentially leading to overlooked vulnerabilities. Additionally, rule-based tools are often limited in scope, typically targeting specific programming languages or certain types of vulnerabilities. LLMs, however, offer a new level of automation and intelligence in this area. LLMs can analyze extensive codebases and detect patterns that are commonly associated with vulnerabilities. These models are trained on massive datasets that include programming languages, software repositories, and vulnerability databases. By pasting the code into an LLM and explicitly requesting it to localize the vulnerability, the LLM can efficiently pinpoint the location of the vulnerable code.
@@
While LLM has demonstrated potential for vulnerability localization, there are several notable limitations. First, LLMs struggle to maintain reasoning accuracy over longer code contexts (Liu et al., 2024b; Kuratov et al., 2024). For instance, in our exploratory experiments, we observed a significant drop in the output accuracy as the lines of code (LoC) increased: when the LoC exceeded 300, the accuracy fell below 5%. Although more advanced reasoning techniques such as Self-Consistency (Wang et al., 2022)-—which analyzes the code multiple times using different reasoning paths to ensure consistency, and rStar (Qi et al., 2024)-—which uses a self-play generation-discrimination process with Monte Carlo Tree Search (MCTS), where two similar models mutually verify reasoning paths—can improve accuracy, these methods still fail to effectively address the issue of long context degradation. Moreover, they rely on stochastic decoding strategies, which introduce randomness into the results and hinder accurate vulnerability localization. Even methods like majority voting resulted in only marginal improvements.
@@
Our work aims to enhance the accuracy of vulnerability localization by leveraging self-attention mechanisms (self-attention allows the model to assign varying importance to different input elements, which is the core of LLMs). The reason lies in the intrinsic connection between self-attention and the model’s output, making it a common tool for interpreting model behavior. We hypothesize that if an LLM can classify vulnerabilities (e.g., types (Yin et al., 2024; Liu et al., 2024a)), it should also inherently possess the capability to pinpoint their exact locations. This is based on the premise that code with vulnerabilities is more likely to significantly influence the model’s attention mechanisms and its subsequent outputs. However, contemporary LLMs face challenges in accurately localizing vulnerabilities due to their processing of extensive contextual information. This often results in key lines or code segments being abstracted into more generalized, higher-level information, causing a loss of detail and clarity. Therefore, the critical task is to refine the model’s focus to concentrate specifically on the lines of code that are vulnerable, enhancing the precision of vulnerability detection and localization.
@@
Building on this idea and our exploratory experiments (refer to §3.2 for validation through a series of experiments), we developed LOVA (LO-cating Vulnerabilities via Attention), with three key stages: code line highlighting, attention calculation, and vulnerability localization. In the code line highlighting stage, two versions of the input are generated: a baseline version and a set of highlighted versions. In the attention calculation stage, attention maps are produced for both versions, and the differences between them are analyzed to assess the impact of highlighting specific lines. Finally, in the vulnerability localization stage, the attention patterns are examined to locate vulnerabilities in specific lines.
@@
The evaluation of LOVA reveals its superiority in localizing vulnerabilities compared to traditional LLM-based approaches across different programming languages and benchmarks. In terms of effectiveness (RQ1), LOVA achieved a significant improvement in F1-scores on both the Big-Vul and CVEFixes-C datasets, outperforming vanilla output and other sophisticated LLM techniques such as CoT, MoA, and rStar by up to 5.3$\times$. This high performance stems from the model’s ability to balance precision and recall, minimizing false positives while maintaining high accuracy in identifying vulnerabilities. Particularly in longer contexts, LOVA demonstrates resilience where other techniques falter, maintaining high Top-N accuracy across various LoC ranges. In terms of scalability (RQ2), LOVA showcases its generalization capabilities by performing effectively across multiple languages, including C, Python, Java, and Solidity. It achieves up to a 14.6$\times$ improvement in contract-level vulnerability localization for smart contracts, underscoring its adaptability to diverse coding environments. The robustness of the system (RQ3) is further demonstrated in its cross-compatibility with various LLM architectures, consistently outperforming baseline methods regardless of the underlying model. Finally, the ablation study (RQ4) highlights the importance of each design component, confirming that the integration of contextual information and attention calculation mechanisms is essential to the method’s overall effectiveness in vulnerability localization.
@@
  • We are the first to discover that the self-attention mechanism of LLMs can be effectively utilized for vulnerability localization. We demonstrate that by tracking changes in attention weights, it is possible to identify specific lines of code that are likely to be vulnerable.
@@ 2.1. LLM Architecture
A large language model (LLM) is an advanced type of artificial intelligence designed to understand and generate human-like text. These models are typically built using deep learning techniques, specifically neural networks with a large number of parameters, trained on vast amounts of text data. LLM is typically based on the transformer architecture (Vaswani, 2017).
@@
A transformer is a type of deep learning model architecture. The key innovation of the transformer architecture is its use of self-attention mechanisms, which allow the model to weigh the importance of different words in a sentence, regardless of their position (which will be introduced next). The transformer architecture consists of an encoder and a decoder: the encoder processes the input sequence into context-rich representations, which the decoder then uses, along with previously generated tokens, to produce the final output sequence. Since both the transformer and its sub-components (the encoder and decoder) are capable of handling sequence-to-sequence generation tasks, and each component has different functions, the development of current LLMs has branched into three main technical routes: encoder-decoder (such as T5 (Raffel et al., 2020)), encoder-only (such as BERT (Devlin, 2018)), and decoder-only (such as GPT (Achiam et al., 2023) and LLaMA (Touvron et al., 2023)). Due to the advantages of higher training efficiency and better inference performance, state-of-the-art models predominantly use the decoder-only architecture. In this paper, we particularly focus on decoder-only LLMs.
2.2. Tokens and Next Token Prediction
@@
Tokens: Tokens serve as the fundamental building blocks that LLMs use to represent and interpret language. Tokenization is a critical step in preparing text for processing by the model. Before handling raw text input, the LLM employs a tokenizer to break the text or code into smaller units known as tokens. These tokens can represent anything from a single character to a whole word, depending on the complexity of the language and the tokenization method. For example, when an LLM processes code, it tokenizes the source code into units such as keywords, identifiers, operators, and literals. Each component of the code is treated as a token, which the model processes individually. Similarly, during the output process, the LLM generates tokens instead of producing full text all at once. The tokenizer then converts these tokens back into readable text or code. The use of tokenization enables LLMs to handle different languages and contexts, regardless of their specific syntax, by converting text or code into tokens that can be systematically processed.
@@
Next Token Prediction: Next token prediction is a core task in training and operating LLM. It involves predicting the next token in a sequence of text based on the preceding tokens. In this process, the model takes an input sequence and generates the most likely next token, continuing this until the desired output length is reached. The training objective of the next token prediction can be formalized by the following loss function:
@@
2.3. Self-Attention Mechanism
@@
In decoder-only LLMs, to maintain the auto-regressive property during the generation process—where each token is generated based solely on the previously generated tokens without access to subsequent tokens—masks are applied during the calculation of attention weights. This ensures that the attention weights for the current token to any subsequent tokens are set to zero. Furthermore, to enable parallel computation of self-attention, the multi-head attention mechanism divides each embedding into multiple parts, each processed by separate attention heads. The results from these heads are then concatenated and linearly transformed to produce the final output. Consequently, this approach allows the model to simultaneously attend to different aspects of the input sequence, thereby enhancing its ability to capture diverse features.
Attention captures the relevance between tokens, serving as an inherent explainability feature of the transformer architecture. There has been significant research (Ghaeini et al., 2018; Lee et al., 2017; Wiegreffe and Pinter, 2019; Serrano and Smith, 2019; Wan et al., 2022; Ma et al., 2024a) on leveraging attention for the interpretability of language models. Previous research (Wan et al., 2022; Ma et al., 2024a) has shown that the self-attention mechanism in LLMs can capture both syntactic and semantic information in code. Specifically, this is achieved by analyzing the attention distribution to identify syntactic or semantic relationships between tokens.
3. Motivation, Key Idea, and Threat Model
@@ 3.1. Motivation
Figure 1. Vulnerable code example
Figure 2. The relationship between Lines of Code (LoC) and Average Top-1 Accuracy, generated by Llama-3.1-8B-Instruct, on all datasets in §5.1
@@
Figure 3. Attention maps for four programs, where (a) represents a vulnerability-free program, (c) shows the program after injecting a vulnerability at line 70 of (a). (b) and (d) represent the programs after highlighting line 70 for (a) and (c), respectively. In each figure, each row corresponds to a line number in the code, while the columns indicate the attention from different decoder layers. Darker colors indicate stronger attention, and the black boxes on the line numbers mark the vulnerable lines identified by the LLM.
While it is true that directly querying LLMs to localize code vulnerabilities, or employing more advanced prompt engineering techniques such as Chain of Thought (Wei et al., 2022; Nong et al., 2024) (i.e., we can apply a step-by-step thought process to systematically break down the potential issues in code) can potentially produce the desired outcome, these methods are constrained by several significant limitations.
First, recent studies (Liu et al., 2024b; Kuratov et al., 2024) have demonstrated that the reasoning capabilities of LLMs deteriorate substantially when working with longer contexts. Since vulnerability localization is inherently a reasoning task that often involves analyzing extended code segments, this degradation poses a significant challenge. For example, for the code shown in Figure 1, when we input the complete file containing the vulnerable code into the LLM for vulnerability localization, only two out of ten attempts successfully identified the vulnerability. Furthermore, as shown in Figure 2, as the lines of code (LoC) increase, the output accuracy of the LLM declines significantly. When the LoC exceeds 300, the accuracy drops to below 5%. While existing advanced reasoning methods (e.g., Self-Consistency (Wang et al., 2022) and rStar (Qi et al., 2024)) can improve reasoning to some extent, they are not fundamentally designed to address long context degradation. Consequently, there is still a significant decline in accuracy when dealing with long contexts.
Second, the advanced reasoning methods mentioned earlier typically employ stochastic decoding strategies to explore different reasoning paths, thereby enhancing accuracy. However, this approach introduces randomness into the output, leading to variability that can result in unstable localization outcomes. Consequently, it becomes challenging to pinpoint the exact line containing the vulnerability. Even when methods like majority voting or scoring are used to aggregate multiple outputs, the overall improvement in accuracy is often marginal (Kang et al., 2024a; Purba et al., 2023).
3.2. Exploratory Experiments and Key Idea